OpenStudy (meilendurcer):

A television set is 36 inches wide and has a diagonal length of 42 inches. To the nearest inch, how tall is the set?

OpenStudy (anonymous):

\[\sqrt{42^2-36^2}\] and a calculator

OpenStudy (anonymous):

it is like the last one \[a^2+36^2=42^2\] solve for \(a^2\) then solve for \(a\) or do it in one step as i wrote above

OpenStudy (meilendurcer):

468

OpenStudy (anonymous):

yes, then take the square root

OpenStudy (meilendurcer):

21?

OpenStudy (anonymous):

\(21.63\) to the nearest 100th